Q3: Evaluate (30 points) i cos? (4 + 2i) What is the principle value of the logarithm (40 points) In(-13 - i)
Answer #1

२.3 i coc curri) il platzisi + 2 enezisi į Lerté-2) + ecuiesz? e2c4i-2) + +2 ési eu = # [excui-a) + ام ] = te +2 - ¿ [6.0183